Given a population where the probability of success is p = 0.40, calculate the probabilities below if a sample of 900 is takeHey! I need some help with these questions! Thanks!
Answer #1

a)

Here, μ = 0.4, σ = 0.0163 and x = 0.39. We need to compute P(X >= 0.39). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (0.39 - 0.4)/0.0163 = -0.61

Therefore,
P(X >= 0.39) = P(z <= (0.39 - 0.4)/0.0163)
= P(z >= -0.61)
= 1 - 0.2709 = 0.7291

b)

Here, μ = 0.4, σ = 0.0163 and x = 0.37. We need to compute P(X <= 0.37). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (0.37 - 0.4)/0.0163 = -1.84

Therefore,
P(X <= 0.37) = P(z <= (0.37 - 0.4)/0.0163)
= P(z <= -1.84)
= 0.0329